ABC Company has set the following standards in order to produce one unit
of its single product:

                    standard quantity           standard price
direct materials     2.7 yards                  ?? per yard
direct labor         2 hours                    $14 per hour
variable overhead    2 hours                    ?? per hour

During August, ABC Company spent $133,770 to purchase direct materials and
had direct labor totaling $189,440. During August, ABC Company used 16,000
yards of direct materials in the production of units. ABC had 2,500 yards
of materials in its August 1 direct material inventory and had 1,200 yards
yards remaining in its direct materials inventory at August 31.

ABC Company reported the following variances for August:

  Direct material price variance ..............  $12,495 unfavorable
  Direct labor rate variance ..................  $24,240 unfavorable
  Total direct labor variance .................  $6,560 favorable
  Variable overhead spending variance .........  $1,900 unfavorable
  Variable overhead efficiency variance .......  $20,900 favorable

Calculate the number of units produced by ABC Company in August.
Answer #1

Direct labor rate variance = (standard rate - actual rate)*actual hours

-24240 = (14 - 189440/actual hours)*actual hours

Actual hours = 11800 hours

Direct labor efficiency variance = total labor variance - labor rate variance

= 6560-(24240) = 30800 favorable

Direct labor efficiency variance = (standard hours - actual hours )*standard rate

30800 = (2*units produced - 11800)*14

Number of units produced = 7000 units